OpenStudy (anonymous):

Tell whether the lines are parallel,perpendicular or neither. 7x-4y=4 , x-4y=3

OpenStudy (anonymous):

perpendicular lines have inverse slope parallel lines have the same slope or a scalar multiple of that slope

OpenStudy (anonymous):

first determine the slopes of the two line. to do that, let's rearrange both equations to slope intercept form.. so we'll have y=7/4x+1 and y=-1/4x-3/4.. so the slope of the 2 lines are 7/4 and -1/4. since both aren't the same or aren't the negative reciprocal of each other, then the lines are neither perp. or par.